# Yamaji effect in models of underdoped cuprates

### v1

@@ III.1 Calculation of resistivity
To compute the $c$-axis resistivity, it is necessary to include a finite interlayer hopping $t_{z}(\mathbf{k})$ between neighboring $\mathrm{CuO}_{2}$ planes. The corresponding Hamiltonian reads
$H_{z}=-\sum_{\mathbf{k},\sigma}2t_{z}(\mathbf{k})\cos(k_{z}c_{\mathrm{lat}})c^{\dagger}_{0;\mathbf{k};\sigma}c_{0;\mathbf{k};\sigma},$ (4)
where $c_{0,\mathbf{k}}$ denotes the physical electron operator that can hop between adjacent layers, and $c_{\mathrm{lat}}$ is the $z$-axis lattice constant. In the SDW theory, $c_{0;\mathbf{k};\sigma}$ coincides with the electron operators appearing in Eq. (1). In contrast, within the ALM theory for the FL* phase, it is generally a linear combination of two types of quasiparticles, $c_{\mathbf{k},\sigma}$ and $\psi_{1;\mathbf{k};\sigma}$, as the physical electron operator $c_{0;\mathbf{k};\sigma}$ is related by a canonical transformation to the operators $c_{\mathbf{k};\sigma}$ and $\psi_{1;\mathbf{k};\sigma}$ in the top two layers. At leading order in $1/J_{\perp}$, where $J_{\perp}$ is the rung exchange coupling between the ancilla layers (see Fig. 1 (c)), this canonical transformation is local, and hence momentum independent (see Appendix A of Ref. [48]). So we can write
@@
where the mixing angle $a$ is small for large $J_{\perp}$; we will treat $a$ as an unknown parameter. Furthermore, one must also consider the possible $\mathbf{k}$ dependence of the interlayer hopping $t_{z}(\mathbf{k})$. For cuprates, it is commonly assumed that $t_{z}(\mathbf{k})$ takes the form $t_{z}(\mathbf{k})=t_{z,0}+t_{z,1}[\cos(k_{x})-\cos(k_{y})]^{2}/4$ , as discussed in Ref. [49]. In what follows, we focus on the case with $t_{z,0}\neq 0$ and defer the discussion of $t_{z,1}$ to a later stage.
Here we use the semiclassical method to compute the $c$-direction conductivity, where the quasiparticles undergo cyclotron motion in a plane perpendicular to the magnetic field. To this end, we use the semi-classical Chamber formula
@@ IV Results
We compute the resistivity as a function of the polar and azimuthal angles $(\theta,\phi)$ of the magnetic field $\mathbf{B}$ for both the FL* and the SDW theory, as shown in Fig. 2. The calculations are performed at a hole doping level $p=0.1$ in $\mathrm{HgBa}_{2}\mathrm{CuO}_{4+\delta}$, using microscopic lattice parameters $a_{\mathrm{lat}}=3.88$$\text{\AA }$ and $c_{\mathrm{lat}}=9.50$$\text{\AA }$. For the bare hopping of electron, we take $t_{c}=0.22$, $t_{c}^{\prime}=-0.034$, $t_{c}^{\prime\prime}=0.036$, $t_{c}^{\prime\prime\prime}=-0.007$, and $t_{z,0}=0.01$ (all in units of eV) for both ALM and SDW theories. In the ALM theory of the FL* phase, we use the ancilla fermion $\psi$ hopping $t_{\psi}=0.1$, $t_{\psi}^{\prime}=-0.03$, $t_{\psi}^{\prime\prime}=-0.01$, and $\Phi=0.09$ [22]. In the SDW, we use $\Delta=0.18$ to fully gap the electron pocket. Results are presented for both a smaller $\omega_{c}\tau$, which better mimics experimental conditions, and a larger $\omega_{c}\tau$, which provides sharper resolution of the Yamaji oscillations.
@@ V Discussion
$\psi_{1}$ Response to EM Field In applying the semi-classical Chambers equation [Eq. (6)] to the ancilla Fermi pocket, we assume that electrons execute cyclotron orbits under the magnetic field. This amounts to assuming that both the $c$ and $\psi_{1}$ fermions respond in the same way to the external electromagnetic field. Such an assumption is justified in two dimensions[21], where condensation of the bosonic field $\Phi$ locks the in-plane external magnetic field to the internal gauge field, $\mathbf{A}_{\parallel}=\mathbf{a}_{\parallel}$, via the Higgs mechanism. In this case, the $\psi_{1}$ particles are effectively coupled to the external field and therefore contribute to transport. In a realistic three-dimensional system with finite interlayer hopping $t_{z}$, whether $\psi_{1}$ couple to the $z$ direction electromagnetic field needs re-justification. We note the finite mixing parameter $a$ in Eq. (5) can induce interlayer correlations of both $c$ and $\psi_{1}$, leading to a finite phase stiffness along the $z$ direction for the bosonic field $\Phi$. The stiffness along $z$ is generally much weaker than the in-plane stiffness, and this anisotropy may introduce additional broadening in the interlayer transport response. We leave it to future to study this effect beyond the current semi-classical calculation.
Momentum dependence of $t_{z}$ We have also examined the effect of $\bf k$ dependene of the interlayer hopping, $t_{z}(\mathbf{k})=t_{z,0}+t_{z,1}[\cos k_{x}-\cos k_{y}]^{2}/4$, which is symmetry-allowed and often relevant in multilayer cuprates. As shown in Appendix C, when we add a $t_{z,1}$ term on top of $t_{z,0}$, the resulting Yamaji angle and overall magnetoresistance profile are nearly identical to those obtained for the isotropic case (Fig. 2). We therefore conclude that the anisotropic component of the interlayer hopping is unimportant in $\mathrm{HgBa}_{2}\mathrm{CuO}_{4+\delta}$ at low hole doping.
@@
ADMR in LSCO Finally, we note that angle-dependent magnetoresistance has also been measured in other cuprate materials, such as LSCO [4], at a higher doping levels ($p\sim 0.21$). In this regime, the magnetic field is not strong enough to induce the Yamaji effect. Instead, the resistivity decreases with increasing $\theta$ at small tilt angles, and develops a peak near $\theta=90^{\circ}$. To qualitatively capture this behavior, we calculate the $c$-axis resistivity of the ancilla pocket assuming an anisotropic interlayer hopping $t_{z}(\mathbf{k})=t_{z,1}[\cos(k_{x})-\cos(k_{y})]^{2}/4$. For smaller $\omega_{c}\tau=0.4$, $\rho_{zz}$ decreases monotonically with $\theta$ and reaches its minimum at $\theta=90^{\circ}$. For larger $\omega_{c}\tau=1.0$, $\rho_{zz}$ first increases to a maximum before decreasing again. The appearance of this maximum signals the onset of a Yamaji-type oscillation, although the first Yamaji angle is smaller than in the isotropic case due to the strong momentum dependence of $t_{z}(\mathbf{k})$. We therefore predict that at sufficiently high magnetic fields, the Yamaji effect should also emerge in LSCO.
On the other hand, the resistivity peak at $\theta=90^{\circ}$ is absent in the present calculation because $v_{z}$ vanishes along the back side of the Fermi pocket, consistent with the findings of Ref. [50]. To reproduce this feature, one needs to include a finite hybridization $a$ between the physical and ancilla layers $c$ and $\psi_{1}$ respectively, as discussed in more detail in Appendix. D.
@@ VI Conclusion
In summary, we have computed the $c$-axis magnetoresistance $\rho_{zz}(\theta,\phi)$ as a function of the angle $(\theta,\phi)$ of the magnetic field within the semiclassical Boltzmann framework for both the SDW and FL* scenarios of the pseudogap phase. Using the ancilla layer model for FL*, we show that the observed Yamaji peak in HgBa2CuO4+δ[5] is quantitatively reproduced by a small Fermi pocket of area $A_{\mathrm{FS}}=p/8$. In contrast, the SDW reconstruction with $A_{\mathrm{FS}}=p/4$ is less consistent with the experiment for the direction $\phi=45^{\circ}$ of the in-plane component of the magnetic field. For example, it predicts an additional Yamaji peak, which is absent in the current experiment.
@@ Acknowledgements.
We thank T. Senthil for valuable discussions. S. S. was supported by the U.S. National Science Foundation grant No. DMR-2245246 and by the Simons Collaboration on Ultra-Quantum Matter which is a grant from the Simons Foundation (651440, S.S.). JYZ and YHZ were supported by the National Science Foundation under Grant No. DMR-2237031.
@@ Appendix A Dependence of Yamaji angle on additional parameters in the FL* phase
$t_{z,\mathrm{eff}}\equiv\frac{v_{z}(\mathbf{k})}{-2t_{z}(\mathbf{k})c_{\mathrm{lat}}\sin(k_{z}c_{\mathrm{lat}})}~,$ (7)
where $v_{z}(\mathbf{k})=\partial\epsilon/\partial k_{z}$ is the out-of-plane velocity of the hybridized band formed by $c$ and $\psi_{1}$, and $2t_{z}(\mathbf{k})c_{\mathrm{lat}}\sin(k_{z}c_{\mathrm{lat}})$ is the bare $z$ velocity computed from the interlayer hopping Eq. (4). For both $a<0$ and $a>0$, the out-of-plane velocity is strongly suppressed on the backside of the Fermi pocket, consistent with the fact that only one side of the pocket carries significant spectral weight $Z_{\mathbf{k}}$. This behavior is qualitatively distinct from the simplified elliptical Fermi pocket used in Ref. [5], where a finite $v_{z}$ is assumed across the entire Fermi surface.
@@
Figure 4: Yamaji angle of ALM for different mixing angles $a$, with $a$ defined in Eq. (5) (a) and (b): the effective $t_{z,\mathrm{eff}}(\mathbf{k})$ as a function of momentum $\mathbf{k}$ for $a=-0.1\pi$ and $a=0.1\pi$. The dashed white line indicate the Fermi pocket and the red start marks the momentum $k^{*}$ of the corner of Fermi pocket. (c) The Yamaji angle for $\phi=0^{\circ}$ as a function of $a$. Data points were obtained from the peak positions of the interlayer resistivity $\rho_{zz}(\theta,0)$ calculated at $\omega_{c}\tau=4.0$. (d) The effective $t_{z,\mathrm{eff}}(\mathbf{k}^{*})$ at the corner of the Fermi pocket $\mathbf{k}^{*}$
Figure 5: Yamaji angle of FL* for different hybridization strength $\Phi$, with $\Phi$ defined in Eq. (2). (a) The Yamaji angle $\theta_{\mathrm{Yamaji}}$ as a function of $\Phi$ for $\phi=0^{\circ}$ and $a=0$, extracted from the date of $\omega_{c}\tau=4.0$. The red star indicate the parameter we used in the main text. (b) Calculated interlayer resistivity $\rho_{zz}(\theta,\phi)/\rho_{zz}(0,0)$ as a function of the polar angle $\theta$ for azimuthal angles $\phi=0^{\circ}$, $22.5^{\circ}$, and $45^{\circ}$, with $\omega_{c}\tau=3.0$.

### v3

@@ III.1 Calculation of resistivity
To compute the $c$-axis resistivity, it is necessary to include a finite interlayer hopping $t_{z}(\mathbf{k}_{2d})$ between neighboring $\mathrm{CuO}_{2}$ planes. The corresponding Hamiltonian reads
$H_{z}=-\sum_{\mathbf{k},\sigma}2t_{z}(\mathbf{k}_{2d})\,\cos(k_{z}c_{\mathrm{lat}})\,c^{\dagger}_{0;\mathbf{k};\sigma}c_{0;\mathbf{k};\sigma},$ (6)
where $c_{0,\mathbf{k}}$ denotes the physical electron operator that can hop between adjacent layers, and $c_{\mathrm{lat}}$ is the $z$-axis lattice constant. Notice that we now have a a full three-dimensional dispersion which depends upon $\mathbf{k}=(\mathbf{k}_{2d},k_{z})$.
In the SDW theory, $c_{0;\mathbf{k};\sigma}$ coincides with the electron operators appearing in Eq. (II). Because we care about the $z$-direction transport, the ordering momentum in $z$ direction now matters. If we assume there is a large correlation length along the $z$-direction, there are two major ansatze with momentum $\mathbf{Q}=(\pi,\pi,\pi)$ and $\mathbf{Q}=(\pi,\pi,0)$, corresponding to antiferromagnetic and ferromagnetic inter-layer spin coupling respectively. We present the results for $\mathbf{Q}=(\pi,\pi,\pi)$ in the appendix, which are very different from the current experiment. In this case the inter-layer hopping is frustrated because of the opposite spin direction between the two nearest-neighbor layers. It turns out that the $z$-direction resistivity $\rho_{zz}$ decreases with $\theta$ at small $\theta$, opposite to the trend in the experiment. In the main text we will mainly focus on the order $\mathbf{Q}=(\pi,\pi,0)$ in the SDW scenario because this may be the best case to match the experiment. However, we note that the correlation length in the $z$-direction is likely smaller than the layer spacing in real system and thus the ordering direction between two layers may be random. In such a scenario, the hopping of hole-pocket quasiparticles between any two layers would depend on the overlap of the in-plane SDW ordering directions of these two layers. The lack of SDW correlations in the $z$-direction can effectively render such hopping random, and this poses a serious challenge to the SDW explanation of the experimental observation of the Yamaji effect [53, 2]. Nevertheless, In the following we will focus on SDW ordering $\mathbf{Q}=(\pi,\pi,0)$ and leave a study of the realistic case with short range SDW order to the future work.
Within the ALM theory for the FL* phase, $c_{0;\mathbf{k};\sigma}$ is generally a linear combination of two types of quasiparticles, $c_{\mathbf{k},\sigma}$ and $\psi_{1;\mathbf{k};\sigma}$, as the physical electron operator $c_{0;\mathbf{k};\sigma}$ is related by a canonical transformation to the operators $c_{\mathbf{k};\sigma}$ and $\psi_{1;\mathbf{k};\sigma}$ in the top two layers. At leading order in $1/J_{\perp}$, where $J_{\perp}$ is the rung exchange coupling between the ancilla layers (see Fig. 1 (c)), this canonical transformation is local, and hence momentum independent (see Appendix A of Ref. [38]). So we can write
@@
where the mixing angle $a$ is small for large $J_{\perp}$; we will mainly use $a=0$ in the main text. Furthermore, one must also consider the possible $\mathbf{k}_{2d}$ dependence of the interlayer hopping $t_{z}(\mathbf{k}_{2d})$. For cuprates, it is commonly assumed that $t_{z}(\mathbf{k}_{2d})$ takes the form
$t_{z}(\mathbf{k}_{2d})=t_{z,0}+t_{z,1}[\cos(k_{x})-\cos(k_{y})]^{2}/4\,$, (8)
as discussed in Ref. [21]. In what follows, we focus on the case with $t_{z,0}\neq 0$ and defer the discussion of $t_{z,1}$ to a later stage.
For both FL* and SDW phases, we use the semiclassical method to compute the $c$-direction conductivity, where the quasiparticles undergo cyclotron motion in a plane perpendicular to the magnetic field. To this end, we use the semi-classical Chamber formula
@@ IV Results
We compute the resistivity as a function of the polar and azimuthal angles $(\theta,\phi)$ of the magnetic field $\mathbf{B}$ for both the FL* and the SDW theory, as shown in Fig. 2. The calculations are performed at a hole doping level $p=0.1$ in $\mathrm{HgBa}_{2}\mathrm{CuO}_{4+\delta}$, using microscopic lattice parameters $a_{\mathrm{lat}}=3.88$$\text{\AA }$ and $c_{\mathrm{lat}}=9.50$$\text{\AA }$. For the bare hopping of electron, we take $t_{c}=0.22$, $t_{c}^{\prime}=-0.034$, $t_{c}^{\prime\prime}=0.036$, $t_{c}^{\prime\prime\prime}=-0.007$, and $t_{z,0}=0.01$ (all in units of eV) for both ALM and SDW theories. In the ALM theory of the FL* phase, we use the ancilla fermion $\psi$ hopping $t_{\psi}=0.1$, $t_{\psi}^{\prime}=-0.03$, $t_{\psi}^{\prime\prime}=-0.01$, and $\Phi=0.09$ [32]. In the SDW, we use $\Delta=0.18$ to fully gap the electron pocket. The effective masses of the hole pockets are estimated to be $1.04m_{e}$ for the ALM theory and $3.44,m_{e}$ for the SDW theory, where $m_{e}$ is the free electron mass. The effective mass is extracted by comparing the cyclotron periods with that of a free hole pocket. While these parameters are not fine-tuned to fit the experimental band structure, the calculated ADMR curve turns out to be stable against changing of the parameters. Results are presented for both a smaller $\omega_{c}\tau$, which better mimics experimental conditions, and a larger $\omega_{c}\tau$, which provides sharper resolution of the Yamaji oscillations.
@@ V Discussion
$\psi_{1}$ Response to EM Field. In applying the semi-classical Chambers equation [Eq. (9)] to the ancilla Fermi pocket, we assume that electrons execute cyclotron orbits under the magnetic field. This amounts to assuming that both the $c$ and $\psi_{1}$ fermions respond in the same way to the external electromagnetic field. Such an assumption is justified in two dimensions[61], where condensation of the bosonic field $\Phi$ locks the in-plane external magnetic field to the internal gauge field, $\mathbf{A}_{\parallel}=\mathbf{a}_{\parallel}$, via the Higgs mechanism. In this case, the $\psi_{1}$ particles are effectively coupled to the external field and therefore contribute to transport. In a realistic three-dimensional system with finite interlayer hopping $t_{z}$, whether $\psi_{1}$ couple to the $z$ direction electromagnetic field needs re-justification. We note the finite mixing parameter $a$ in Eq. (7) can induce interlayer correlations of both $c$ and $\psi_{1}$, leading to a finite phase stiffness along the $z$ direction for the bosonic field $\Phi$. The stiffness along $z$ is generally much weaker than the in-plane stiffness, and this anisotropy may introduce additional broadening in the interlayer transport response. We leave it to future to study this effect beyond the current semi-classical calculation.
Momentum dependence of $t_{z}$. We have also examined the effect of $\bf k$ dependene of the interlayer hopping $t_{z}(\mathbf{k}_{2d})$ in Eq. (8), which is symmetry-allowed and often relevant in multilayer cuprates. As shown in Appendix C, when we add a $t_{z,1}$ term on top of $t_{z,0}$, the resulting Yamaji angle and overall magnetoresistance profile are nearly identical to those obtained for the isotropic case (Fig. 2). We therefore conclude that the anisotropic component of the interlayer hopping is unimportant in $\mathrm{HgBa}_{2}\mathrm{CuO}_{4+\delta}$ at low hole doping.
@@
Short range SDW. In the main text, we assume the spin-density-wave (SDW) has a long correlation length. Even in this idealized case, we have shown that it is in poorer agreement with the experiment than the FL* phase. In reality, the correlation length of the SDW order is known to be short in certain cuprates, like HgBa2CuO4+δ [53, 2]. For instance, in the $z$-direction, the relative Néel ordering between adjacent layers may be uncorrelated, which would effectively randomize the inter-layer hopping. We expect the agreement with the experiment to be even poorer in this scenario, but future work is needed to rigorously model the short-range SDW case.
ADMR in LSCO. Finally, we note that angle-dependent magnetoresistance has also been measured in other cuprate materials, such as LSCO [12], at a higher doping levels ($p\sim 0.21$). In this regime, the magnetic field is not strong enough to induce the Yamaji effect. Instead, the resistivity decreases with increasing $\theta$ at small tilt angles, and develops a peak near $\theta=90^{\circ}$. To qualitatively capture this behavior, we calculate the $c$-axis resistivity of the ancilla pocket assuming an anisotropic interlayer hopping in Eq. (15). For smaller $\omega_{c}\tau=0.4$, $\rho_{zz}$ decreases monotonically with $\theta$ and reaches its minimum at $\theta=90^{\circ}$. For larger $\omega_{c}\tau=1.0$, $\rho_{zz}$ first increases to a maximum before decreasing again. The appearance of this maximum signals the onset of a Yamaji-type oscillation, although the first Yamaji angle is smaller than in the isotropic case due to the strong momentum dependence of $t_{z}(\mathbf{k}_{2d})$. We therefore predict that at sufficiently high magnetic fields, the Yamaji effect should also emerge in LSCO.
On the other hand, the resistivity peak at $\theta=90^{\circ}$ is absent in the present calculation because $v_{z}$ vanishes along the back side of the Fermi pocket, consistent with the findings of Ref. [36]. To reproduce this feature, one needs to include a finite hybridization $a$ between the physical and ancilla layers $c$ and $\psi_{1}$ respectively, as discussed in more detail in Appendix. F.
Currently, the Yamaji effect has been reported only in Hg-based cuprates, ikely due to their minimal disorder. It is possible that the Yamaji effect could be observed in other materials under stronger magnetic fields $B$. Note that DFT calculations have reported conduction bands derived from Hg-O chains lying above the Fermi level. However, these bands, with reported energies of the order of $0.2$ eV – 1 eV[47, 35], are not accessible with current magnetic field strengths.
@@ VI Conclusion
In summary, we have computed the $c$-axis magnetoresistance $\rho_{zz}(\theta,\phi)$ as a function of the angle $(\theta,\phi)$ of the magnetic field within the semiclassical Boltzmann framework for both the SDW and FL* scenarios of the pseudogap phase. Using the ancilla layer model for FL*, we show that the observed Yamaji peak in HgBa2CuO4+δ[6] is quantitatively reproduced by a small Fermi pocket of area $A_{\mathrm{FS}}=p/8$. In contrast, the SDW scenario highly depends on the ordering in the $z$ direction. SDW with momentum $\mathbf{Q}=(\pi,\pi,\pi)$ qualitatively disagree with the current experiment. For SDW with momentum $\mathbf{Q}=(\pi,\pi,0)$ gives a reconstruction with $A_{\mathrm{FS}}=p/4$ and is less consistent with the experiment for the direction $\phi=45^{\circ}$ of the in-plane component of the magnetic field. For example, it predicts an additional Yamaji peak, which is absent in the current experiment. We conclude that FL* phase better explains the current experimental observation at $p=0.1$ and we propose future experiments to study the full doping dependence to verify the Fermi surface area $A_{\mathrm{FS}}={p}/{8}$.
@@ Acknowledgements.
We thank Mun Chan, A. Chubukov, M. Greven, N. Harrison, C.M. Jian, J. Kim, P.A. Lee, A. Nikolaenko, P. Nosov, B. Ramshaw, T. Senthil, and L. Taillefer for valuable discussions. SS was supported by the U.S. National Science Foundation grant No. DMR-2245246 and by the Simons Collaboration on Ultra-Quantum Matter which is a grant from the Simons Foundation (651440, S.S.). JYZ and YHZ were supported by the National Science Foundation under Grant No. DMR-2237031.
@@ Appendix A Dependence of Yamaji angle on additional parameters in the FL* phase
$t_{z,\mathrm{eff}}(\mathbf{k})\equiv\frac{v_{z}(\mathbf{k})}{-2t_{z}(\mathbf{k}_{2d})c_{\mathrm{lat}}\sin(k_{z}c_{\mathrm{lat}})}~,$ (10)
where $v_{z}(\mathbf{k})=\partial\epsilon/\partial k_{z}$ is the out-of-plane velocity of the hybridized band formed by $c$ and $\psi_{1}$, and $2t_{z}(\mathbf{k}_{2d})c_{\mathrm{lat}}\sin(k_{z}c_{\mathrm{lat}})$ is the bare $z$ velocity computed from the interlayer hopping Eq. (6). In the limit of small $t_{z}$, the $k_{z}$-dependence of $v_{z}(\mathbf{k})/\sin(k_{z}c_{\mathrm{lat}})$ is negligible, allowing the approximation $t_{z,\mathrm{eff}}(\mathbf{k})\approx t_{z,\mathrm{eff}}(\mathbf{k}_{2d})$. For both $a<0$ and $a>0$, the out-of-plane velocity is strongly suppressed on the backside of the Fermi pocket, consistent with the fact that only one side of the pocket carries significant spectral weight $Z_{\mathbf{k}}$. This behavior is qualitatively distinct from the simplified elliptical Fermi pocket used in Ref. [6], where a finite $v_{z}$ is assumed across the entire Fermi surface.
@@
Figure 4: Yamaji angle of ALM for different mixing angles $a$, with $a$ defined in Eq. (7) (a) and (b): the effective $t_{z,\mathrm{eff}}(\mathbf{k}_{2d})$ as a function of momentum $\mathbf{k}$ for $a=-0.1\pi$ and $a=0.1\pi$. The dashed white line indicate the Fermi pocket and the red start marks the momentum $k^{*}$ of the corner of Fermi pocket. (c) The Yamaji angle for $\phi=0^{\circ}$ as a function of $a$. Data points were obtained from the peak positions of the interlayer resistivity $\rho_{zz}(\theta,0)$ calculated at $\omega_{c}\tau=4.0$. (d) The effective $t_{z,\mathrm{eff}}(\mathbf{k}_{2d}^{*})$ at the corner of the Fermi pocket $\mathbf{k}_{2d}^{*}$
Figure 5: Yamaji angle of FL* and SDW for different hybridization strength $\Phi$ and $\Delta$, with $\Phi$ defined in Eq. (4) and $\Delta$ defined in (II). (a) The Yamaji angle of FL* $\theta_{\mathrm{Yamaji}}$ as a function of $\Phi$ for $\phi=0^{\circ}$ and $a=0$, extracted from the date of $\omega_{c}\tau=4.0$. The red star indicate the parameter we used in the main text. (b) Calculated interlayer resistivity $\rho_{zz}(\theta,\phi)/\rho_{zz}(0,0)$ for $\Phi=0.15$ as a function of the polar angle $\theta$ for azimuthal angles $\phi=0^{\circ}$, $22.5^{\circ}$, and $45^{\circ}$, with $\omega_{c}\tau=3.0$. (c) The Yamaji angle of SDW $\theta_{\mathrm{Yamaji}}$ as a function of $\Delta$ for $\phi=0^{\circ}$, extracted from the date of $\omega_{c}\tau=4.0$. The red star indicate the parameter we used in the main text. (d) Calculated interlayer resistivity $\rho_{zz}(\theta,\phi)/\rho_{zz}(0,0)$ for $\Delta=0.30$ as a function of the polar angle $\theta$ for azimuthal angles $\phi=0^{\circ}$, $22.5^{\circ}$, and $45^{\circ}$, with $\omega_{c}\tau=3.0$.